A body of mass m is released from top of a rough inclined plane, if the force of friction is f,then the body Will reach the bottom with what velocity ?

Grade:12th pass

1 Answers

Piyush Kumar Behera
7 years ago
@Parul Rajput
See in this type of problem se work energy principle,which states that,
the total work done on the body=change in kinetic energy.
 
In this question initial kinetic energy=0
Let final kinetic energy be K,
length of incline=hieght x cosec\theta where \theta is the angle of inclination
using work-energy theorem,
 
work done by gravity+work done by friction=K-0
mg x hieght of the wedge – f x length of the inclined plane= K
mg x hieght of the wedge – f x hieght x cosec\theta=1/2mv2
solve further to get the answer..
Hope that helps!!!  please approve to encourage me 
thank you